# Practitioners’ Challenges and Perceptions of CI Build Failure Predictions at Atlassian

### v1

@@ Abstract.
conference: ; July 15–19, 2024; Brazilcopyright: none
@@ 1. Introduction
Despite these advantages, CI comes with its own challenges. At Atlassian  [^1] https://www.atlassian.com/, one of the challenges is to reduce CI build failures on the master branch. CI build failures on the master branch could significantly impact the software development process and development teams. For example, they could delay rollouts of new features or security patches, require rollbacks to fix the issues, and hamper the productivity of software engineers (Ananthanarayanan et al., 2019). From 2021 to 2023, we see that 17% of the CI builds on the master branch failed in Atlassian’s projects, resulting in an average of 10 days of wasted time per project. Such disruptions not only prevent developers from proceeding further with the development but also incur a high cost for organizations to rectify. Therefore, this challenge necessitates a thorough investigation of the CI build failures at Atlassian.
To mitigate CI build failures, various CI build prediction techniques (Chen et al., 2020; Saidani et al., 2020; Santos et al., 2022; Kawalerowicz and Madeyski, 2023; Saidani et al., 2022; Xia and Li, 2017) have been developed to preemptively detect states in the CI process that are likely to cause build failures. This enables developers to take the necessary actions and avoid these failures. However, despite the availability of these techniques, there still remain gaps between industry and research. First, prior studies are based on open-source projects, which may not generalize to the industrial context with different scales. In addition, while explainability is an important aspect of their adoption by developers, little is known how developers perceive about the CI build failure predictions and their explanations.
In our study, we conducted empirical studies to investigate the factors that are associated with CI build failures and practitioners’ perceptions of utilizing CI build predictions at Atlassian. Specifically, we analyzed 350,037 pull requests from 1,630 projects at Atlassian and investigated 11 factors that are possibly associated with CI build failures. We then conducted a survey study with Atlassian developers to gain a better understanding of the adoption of CI build failure predictions. We investigated four research questions:
@@ 1. RQ1)
    We performed a quantitative analysis to identify the build-failure-related factors that are pertinent in the context of commercial projects at Atlassian. We found that factors in the repository dimension (e.g., the ratio of the builds that failed out of the recent five builds of the repository and the ratio of all historical builds of the repository that failed) play the most important role in influencing CI build outcomes.
@@ 2. RQ2)
    From the developers’ perspective, while CI build failures could impact the development process and productivity of the team, they could also provide the opportunity for developers to learn from build failures and avoid more serious issues in the future.
@@ 3. RQ3)
    Atlassian developers perceived that CI build predictions could provide proactive insight and aid in the team’s decision-making. However, they noted several challenges that need to be addressed before adopting the predictions in practice, such as the potential for developers’ over-reliance on them.
@@ 4. RQ4)
    Atlassian developers agreed with explanations and suggestions based on the percentage of current changed files that were involved in the previous failed builds and the number of changed files. Developers point out that while the explanations and suggestions for factors related to CI build failures could help developers understand CI build failure, they should be optimized to fit the context of development scenarios.
The paper is organized as follows. Section 2 introduce the CI build process at Atlassian. Section 3 introduces the research methodology of our study. Section 4 presents the results. Section 5 discusses the implications to researchers and practitioners. Section 6 provides the threats to the validity. Section 7 reviews related work. Section 8 draws the conclusion.
@@ 2. The CI build process at Atlassian
However, the CI build of the PR may fail even if it is carefully reviewed. The build failure could significantly impact the software development process at Atlassian. Therefore, prior studies proposed various CI build prediction techniques to early predict the build outcomes (see Figure 1). If the predictions can early indicate that the build will fail, developers can examine the PR again before sending it to the CI process. This enables developers to take the necessary actions to avoid potential CI build failures. If the predictions indicate that the build will succeed, the PR will be delivered to the CI build process and merged into the master branch.
In this work, the close integration of Atlassian’s development teams with Bitbucket gives the opportunity to investigate the CI build processes in situ and gain insights into developers’ perceptions and expectations if integrating the CI build prediction into Bitbucket.
@@ 3.1. Research Questions
We present the motivation of our research questions below:
@@ 3.2. Studied Dataset
To address our research questions, we conducted an empirical study on software projects created by development teams at Atlassian. These projects employ a workflow that streamlines the integration of new code changes (see Section 2). The data relating to this workflow is collected by Atlassian developers and stored in a database for in-depth data analysis.
@@
Data Cleaning.  After selecting the projects, we cleaned the data to ensure the accuracy of the study results. We first excluded the builds that are not triggered on the master branch. Then, we performed the following steps for data cleaning.
@@
Dataset.  We collected the data of PRs merged into the master branch between January 1, 2021, and September 1, 2023. After applying the aforementioned inclusion/exclusion criteria to the data, our dataset consists of a total of 350,037 PRs from 1,630 studied projects.
@@
Table 2. Survey questions (excluding demographics questions)
@@ 3.3. Explanatory Factors
Note that the scope of data available for our study is constrained. First, the data only contains PR level, repository level, and team-related information. Secondly, to strictly comply with Atlassian’s privacy policies, we are requested to exclude some parts of the data (i.e., personally identifiable data and user-generated contents) to ensure the protection of sensitive information.
@@ 3.4. Data Analysis
Similarly to past studies (Thongtanunam et al., 2017; Tantithamthavorn and Hassan, 2018), we aim to understand the effect of the selected factors on the CI build outcome of a PR. To this end, we fit a Logistic Regression Model (LRM). LRM models are commonly used to capture the relationship between the explanatory variables (i.e., factors described in Table 1) and a response variable (i.e., the CI build outcome). In our study, the response variable is assigned the value of TRUE if a build fails, and FALSE otherwise.
@@ 3.5. Survey Design and Participants
Respondent Demographics.  The respondents are software engineers (91%) and software engineering managers (7%). For the experience in the IT industry of respondents , 33% of participants had more than five years of experience, 59% of respondents had one to five years, and 7% of respondents had less than one year. In terms of the frequency of submitting PRs, 46% of respondents have submitted PR more than once a day; 41% of the respondents submitted PR once a day. The remaining respondents submitted PR once a week. Regarding the frequency of respondents experiencing CI build failures, 4% of respondents experienced failures in more than 75% of PRs, 43% in 50%-75% of PRs, 37% in 25%-50% of PRs, and the remaining respondents experienced failures in less than 25% of PRs.
@@ 3.6. Card Sorting
For the analysis of open-ended survey responses, we employ a thematic analysis approach to ensure our analysis remains unbiased and free from preconceptions about the responses. Therefore, similar to previous works (Kononenko et al., 2018, 2016), we used an open coding approach to identify themes and group the data. The first author split the 53 survey responses into isolated quotes (cards). Each card represents a single statement distinct from other statements in a particular response to a question. After that, two of the other authors served as coders to manually label cards. They discuss any uncertainties in the cards to share a common understanding of the data. Finally, similar cards are merged into common themes for further analysis.
@@ 4. Results
4.1. RQ1: What are the most important factors associated with CI Build Failures at Atlassian?
@@
$\chi^{2}$ $p$-value Shifted Value Odds Ratio Wald $\chi^{2}$ 45,235 *** PR Dimension changes_num 0.1% *** 1.0→8.0 0.5%↑ comment_num 0.1% *** 0.0→4.0 1.6%↑ reviewer_num 0.0% ** 2.0→7.0 2.2%↓ wait_time_to_review 0.0% *** 0.0→13.0 0.3%↑ reviewing_time 0.0% ** 0.0→22.0 0.3%↑ per_failed_file 1.4% *** 0.0→0.8 31.6%↑ Repository Dimension repo_prev_build 6.1% *** Passed→Failed 174.7%↑ repo_rec_build 21.4% *** 0.0→0.2 111.0%↑ repo_hist_build 5.6% *** 0.0→0.2 29.6%↑ Contributor Dimension team_num 0.2% *** 1.0→2.0 5.8%↑ team_member_num 0.1% *** 7.0→15.0 2.0%↓ Statistical significance of explanatory power according to Wald $\chi^{2}$ likelihood ratio test: ${}^{\circ}p\geq 0.05;^{*}0.01<p<0.05;^{**}0.001<p<0.01;^{***}p<0.001$:This explanatory variable has a positive relationship with the response variable. ↓:This explanatory variable has a negative relationship with the response variable.
Figure 2. The relationship between the explanatory variable (x-axis) and the likelihood of a build failing (y-axis). The larger the odds value is, the higher the likelihood that the build failing in CI. The gray area represents the 95% confidence interval.
@@
The ratio of the builds that failed out of the recent five builds of the repository shares a positive relationship with the likelihood of CI build failure. Table 4 also shows that the ratio of the builds that failed out of the recent five builds in the repository shares a significant relationship with the likelihood that a build will fail. The likelihood increases by 111% when the ratio value increases from 0 to 0.2. Figure 2 also shows an increasing trend in the likelihood of CI build failure as the value of this variable increases. The narrowing of the confidence interval (gray area) suggests a high precision of this variable to determine the CI build failure. This result indicates that a build tends to fail if many recent failures occured in the CI builds.
@@
4.2. RQ2: What are practitioners’ perceptions about CI build failures?
To answer RQ2, we asked three questions in our survey to capture insights from experienced developers who actively performed CI builds. Q1 is a five-point Likert question that aims to quantify the level of challenges faced by developers when resolving the CI build failures. Q2 is an open-ended question to ask the impact of CI build failures. Q3 is a multiple-selection question with optional free-text responses to gather developers’ opinions about the causes behind CI build failures occurring during the CI builds triggered on the master branch.
@@ 4.2. RQ2: What are practitioners’ perceptions about CI build failures?
RQ2: Atlassian’s developers perceived that CI build failures are challenging to resolve. The responses from developers noted that CI build failures delay the software development process and reduce team productivity. But they also provide opportunities for developers to take the lessons from failing and avoid later more serious issues.
4.3. RQ3: What are practitioners’ perceptions on the usefulness of CI build prediction?
@@
Factors Explanations Suggestions
per_failed_file
High percentage of changed files that were involved in the
previous failed build may cause the previous issues again.
Ensure all files in previous failed build
are carefully checked.
changes_num
A PR with a large number of changed files may contain
more issues.
Break the large PR down into smaller PRs,
each addressing small and coherent changes.
comment_num
A high number of comments may suggest more potential
issues or disagreements in the PR.
Ensure all comments are addressed.
@@ A PR with a large number of changed files may contain
RQ3: Atlassian practitioners acknowledged the value of the CI build prediction in providing proactive insight into PRs and facilitating decision-making. However, the practitioners still had concerns about the accuracy, the risk of over-reliance, and providing balanced and valuable support to all developers across different skill levels.
4.4. RQ4: What are practitioners’ perceptions on the explanations and suggestions of CI build predictions?
@@
RQ4: Atlassian developers agreed that CI build failures are associated with the percentage of current changed files that were involved in the previous failed builds and the number of changed files. However, while recognizing the significance of these factors in understanding the CI build failure, developers emphasized that the explanations and suggestions for these factors should be optimized to fit the context of development scenarios. This underscores a need for further research and investigation to refine and tailor these insights for varied development scenarios.
@@ 5. Implications
Below, we discuss the implications of our work to researchers and practitioners.
@@ 5.1. Suggestions for Researchers
Enhancement of Build Prediction Explainability: Our observation pinpoints the importance of explainability in CI build prediction. We found that developers are more likely to trust and effectively utilize these techniques when they understand the rationale behind predictions. Although we give a glance into the explanation of the CI build prediction, there’s substantial room for advancement. Future research can focus on creating intuitive methods to convey complex model decisions such as interactive visualizations or clear narrative explanations. Studies could also explore the impact of different explanation types on developer trust and model utilization, aiming to tailor explainability features to developers’ specific needs and preferences. Enhancing explainability plays an important role in fostering relationships between developers and CI build prediction.
Investigation into Contextual Dependencies: Our findings highlight that the effectiveness of CI build predictions depends on the context, such as project characteristics and the individual experience levels of developers. This observation underscores the potential benefits of tailoring the level of automation and the nature of interactions between the build prediction technique and its users based on these contexts. It suggests that a one-size-fits-all approach may not be optimal and that customization could enhance the build prediction’s effectiveness and user satisfaction. It is worth conducting further research involving in-depth case studies or comparative analyses across diverse contexts to understand the contextual factors affecting the utilization of CI builds prediction in practice.
Collaboration between Human Judgment and Automated Predictions: The findings in our work point to the necessity of a balanced collaboration between human judgment and automated predictions of build outcomes in CI processes. There’s a need for studies that explore the dynamics of trust, reliance, and decision-making between developers and CI build prediction. For example, researchers can optimize the integration of these techniques by studying how developers interact with and interpret the predictions and suggestions provided. Such research could involve developing frameworks that facilitate effective communication between the CI build prediction technique and the developer, ensuring that the predictions are not only accurate but also presented in a manner that is aligned with the developer’s cognitive and decision-making processes.
5.2. Implication for Practitioners

### v2

@@ Abstract.
Continuous Integration, CI/CD, DevOps, CI Build Failure
ccs: Software and its engineeringcopyright: acmlicenseddoi: 10.1145/3663529.3663856journalyear: 2024submissionid: fsecomp24industry-p93-pisbn: 979-8-4007-0658-5/24/07conference: Companion Proceedings of the 32nd ACM International Conference on the Foundations of Software Engineering; July 15–19, 2024; Porto de Galinhas, Brazilbooktitle: Companion Proceedings of the 32nd ACM International Conference on the Foundations of Software Engineering (FSE Companion ’24), July 15–19, 2024, Porto de Galinhas, Brazil
@@ 1. Introduction
Despite these advantages, CI comes with its own challenges. At Atlassian, one of the challenges is to reduce CI build failures on the main branch. CI build failures on the main branch could significantly impact the software development process and development teams. For example, they could delay rollouts of new features or security patches, require rollbacks to fix the issues, and hamper the productivity of software engineers (Ananthanarayanan et al., 2019). From 2021 to 2023, we observed that failed CI builds on the main branch resulted in an average of 120 hours of wasted build time per project per year in the studied Atlassian’s projects. Such disruptions not only prevent developers from proceeding further with the development but also incur a high cost for organizations to rectify. Therefore, this challenge necessitates a thorough investigation of the CI build failures at Atlassian.
To mitigate CI build failures, various CI build prediction techniques (Chen et al., 2020; Saidani et al., 2020; Santos et al., 2022; Kawalerowicz and Madeyski, 2023; Saidani et al., 2022; Xia and Li, 2017) have been developed to preemptively detect states in the CI process that are likely to cause build failures. This enables developers to take the necessary actions and avoid these failures. However, despite the availability of these techniques, there still remain gaps between industry and research. First, prior studies are based on open-source projects, which may not generalize to the industrial context with different scales. In addition, while explainability is an important aspect of their adoption by developers (Tantithamthavorn and Jiarpakdee, 2021; Jiarpakdee et al., 2020, 2021; Tantithamthavorn et al., 2015; Pornprasit et al., 2021; Tantithamthavorn et al., 2020), little is known how developers perceive about the CI build failure predictions and their explanations.
In our study, we conducted empirical studies to investigate the factors that are associated with CI build failures and practitioners’ perceptions of utilizing CI build predictions at Atlassian. Specifically, we analyzed 350,037 pull requests from 1,630 projects at Atlassian and investigated 11 factors that are possibly associated with CI build failures. We then conducted a survey study to gain a better understanding of the adoption of CI build failure predictions. The survey had 53 respondents (i.e., Atlassian practitioners who participated). We answered four research questions below:
@@ 1. RQ1)
    We found that factors in the repository dimension (e.g., the ratio of the builds that failed out of the recent five builds of the repository and the ratio of all historical builds of the repository that failed) play the most important role in influencing CI build outcomes.
@@ 2. RQ2)
    From the respondents’ perspective, while CI build failures could impact the development process and productivity of the team, they could also provide the opportunity for developers to learn from build failures and avoid more serious issues in the future.
@@ 3. RQ3)
    Respondents perceived that CI build predictions could provide proactive insight and aid in the team’s decision-making. However, they noted several challenges that need to be addressed before adopting the predictions in practice, such as the potential for developers’ over-reliance on them.
@@ 4. RQ4)
    Respondents agreed with explanations and suggestions based on the percentage of current changed files that were involved in the previous failed builds and the number of changed files. Developers pointed out that the explanations and suggestions for the related factors could help developers understand CI build failure. However, they also mentioned that the explanations and suggestions should be optimized to fit the context of development scenarios.
Our study sheds light on the challenges and expectations involved in integrating CI build prediction tools, providing valuable insights for enhancing CI processes.
@@ 2. The CI build process at Atlassian
However, the CI build of the PR may fail even if it is carefully reviewed. The build failure could significantly impact the software development process at Atlassian. To mitigate such a problem, prior studies proposed various CI build prediction techniques to early predict the build outcomes (see Figure 1). If the predictions can early indicate that the build will fail, developers can examine the PR again before sending it to the CI process. This enables developers to take the necessary actions to avoid potential CI build failures. If the predictions indicate that the build will succeed, the PR will be delivered to the CI build process and merged into the main branch of the code repository.
In this work, the close integration of Atlassian’s development teams with Bitbucket gives the opportunity to investigate the CI build processes in situ and gain insights into developers’ perceptions and expectations if integrating CI build predictions into Bitbucket.
@@ 3.1. Research Questions
We present the motivation for our research questions below:
@@ 3.2. Studied Dataset
To address our research questions, we conducted an empirical study on Atlassian’s internal software projects. These projects are created by Atlassian development teams (not Atlassian’s customers). In these projects, Atlassian developers employ a workflow that streamlines the integration of new code changes (see Section 2). The data relating to this workflow is collected by Atlassian teams and stored in a database for in-depth data analysis.
@@
Data Cleaning.  After selecting the projects, we cleaned the data to ensure the accuracy of the study results. We first excluded the builds that are not triggered on the main branch. Then, we performed the following steps for data cleaning.
@@
Dataset.  We collected the data of PRs merged into the main branch between January 1, 2021, and September 1, 2023. After applying the aforementioned inclusion/exclusion criteria to the data, our dataset consists of a total of 350,037 PRs from 1,630 Atlassian projects.
@@
Table 2. Survey questions (excluding demographics questions).
@@ 3.3. Explanatory Factors
Note that the scope of data available for our study is constrained. First, the data only contains PR level, repository level, and team-related information. Secondly, to strictly comply with Atlassian’s privacy policies, we are required to exclude any data, including but not limited to personally identifiable data and user-generated content to ensure the protection of sensitive information.
@@ 3.4. Data Analysis
Similarly to past studies (Thongtanunam et al., 2017; Tantithamthavorn and Hassan, 2018), we aim to understand the effect of the selected factors on the CI build outcome of a PR at Atlassian. To this end, we fit a Logistic Regression Model (LRM). LRM models are commonly used to capture the relationship between the explanatory variables (i.e., factors described in Table 1) and a response variable (i.e., the CI build outcome). In our study, the response variable is assigned the value of TRUE if a build fails, and FALSE otherwise.
@@ 3.5. Survey Design and Participants
Respondent Demographics.  The respondents are software engineers (91%) and software engineering managers (7%). For the experience in the IT industry of respondents , 33% of participants had more than five years of experience, 59% of respondents had one to five years, and 7% of respondents had less than one year. In terms of the frequency of submitting PRs, 46% of respondents have submitted PR more than once a day; 41% of the respondents submitted PR once a day. The remaining respondents submitted PR once a week. Regarding the frequency of respondents experiencing CI build failures, 4% of respondents experienced failures in more than 75% of PRs, 43% in 50%-75% of PRs, 37% in 25%-50% of PRs, and 16% of respondents experienced failures in less than 25% of PRs.
@@ 3.6. Card Sorting
For the analysis of open-ended survey responses, we employ a thematic analysis approach to ensure our analysis remains unbiased and free from preconceptions about the responses. Therefore, similar to previous works (Kononenko et al., 2018, 2016), we used an open coding approach to identify themes and group the data. The first author split the 53 survey responses into isolated quotes (cards). Each card represents a single statement distinct from other statements in a particular response to a question. After that, two of the other authors independently served as coders to manually label cards. They discuss any uncertainties in the cards to share a common understanding of the data. Finally, similar cards are merged into common themes for further analysis.
@@ 4. Results
RQ1: What are the most important factors associated with CI Build Failures at Atlassian?
@@
$\chi^{2}$ $p$-value Shifted Value Odds Ratio Wald $\chi^{2}$ 45,235 *** PR Dimension changes_num 0.1% *** 1.0→8.0 0.5%↑ comment_num 0.1% *** 0.0→4.0 1.6%↑ reviewer_num 0.0% ** 2.0→7.0 2.2%↓ wait_time_to_review 0.0% *** 0.0→13.0 0.3%↑ reviewing_time 0.0% ** 0.0→22.0 0.3%↑ per_failed_file 1.4% *** 0.0→0.8 31.6%↑ Repository Dimension repo_prev_build 6.1% *** Passed→Failed 174.7%↑ repo_rec_build 21.4% *** 0.0→0.2 111.0%↑ repo_hist_build 5.6% *** 0.0→0.2 29.6%↑ Contributor Dimension team_num 0.2% *** 1.0→2.0 5.8%↑ team_member_num 0.1% *** 7.0→15.0 2.0%↓ Statistical significance of explanatory power according to Wald $\chi^{2}$ likelihood ratio test: ${}^{\circ}p\geq 0.05;^{*}0.01<p<0.05;^{**}0.001<p<0.01;^{***}p<0.001$: The explanatory variable has a positive relationship with the response variable. ↓: The explanatory variable has a negative relationship with the response variable.
Figure 2. The relationship between the explanatory variable (x-axis) and the likelihood of a build failing (y-axis). The larger the odds value, the higher the likelihood that a build will fail. The gray area represents the 95% confidence interval.
@@
The ratio of the builds that failed out of the recent five builds of the repository shares a positive relationship with the likelihood of CI build failure. Table 4 also shows that the ratio of the builds that failed out of the recent five builds in the repository shares a significant relationship with the likelihood that a build will fail. The likelihood increases by 111% when the ratio value increases from 0 to 0.2. Figure 2 also shows an increasing trend in the likelihood of CI build failure as the value of this variable increases. The narrowing of the confidence interval (gray area) suggests a high precision of this variable to determine the CI build failure. This result indicates that a build tends to fail if many recent failures occur in the CI builds.
@@
RQ2: What are practitioners’ perceptions about CI build failures?
To answer RQ2, we asked three questions in our survey to capture insights from experienced developers who actively performed CI builds. Q1 is a five-point Likert question that aims to quantify the level of challenges faced by developers when resolving the CI build failures. Q2 is an open-ended question to ask the impact of CI build failures. Q3 is a multiple-selection question with optional free-text responses to gather developers’ opinions about the causes behind CI build failures occurring during the CI builds triggered on the main branch.
@@
RQ2: Respondents perceived that CI build failures are challenging to resolve. The responses from developers noted that CI build failures delay the software development process and reduce team productivity. But they also provide opportunities for developers to take the lessons from failing and avoid later more serious issues.
RQ3: What are practitioners’ perceptions on the usefulness of CI build prediction?
@@
Factors Explanations Suggestions per_failed_file High percentage of changed files that were involved in the previous failed build may cause the previous issues again. Ensure all files in previous failed build are carefully checked. changes_num A PR with a large number of changed files may contain more issues. Break the large PR down into smaller PRs, each addressing small and coherent changes. comment_num A high number of comments may suggest more potential issues or disagreements in the PR. Ensure all comments are addressed.
@@
RQ3: Respondents acknowledged the value of the CI build prediction in providing proactive insight into PRs and facilitating decision-making. However, the practitioners still had concerns about the accuracy, the risk of over-reliance, and providing balanced and valuable support to all developers across different skill levels.
RQ4: What are practitioners’ perceptions on the explanations and suggestions of CI build predictions?
@@
RQ4: Respondents agreed that CI build failures are associated with the percentage of current changed files that were involved in the previous failed builds and the number of changed files. However, while recognizing the significance of these factors in understanding the CI build failure, developers emphasized that the explanations and suggestions for these factors should be optimized to fit the context of development scenarios. This underscores a need for further research and investigation to refine and tailor these insights for varied scenarios.
@@ 5. Implications
Below, we provide the discussion about the implications of our work to researchers and practitioners.
@@ 5.1. Suggestions for Researchers
Collaboration between Human Judgment and Automated Predictions: RQ3 shows the necessity of a balanced collaboration between human judgment and automated predictions of build outcomes in CI processes. There’s a need for detailed studies that explore the dynamics of trust, reliance, and decision-making between developers and CI build prediction. For example, researchers can optimize the integration of these techniques by studying how developers interact with and interpret the predictions and suggestions provided. Such research could involve developing frameworks that facilitate effective communication between the CI build prediction technique and the developer, ensuring that the predictions are not only accurate but also presented in a manner that is aligned with the developer’s cognitive and decision-making processes.
Investigation into Contextual Dependencies: RQ4 highlights that the effectiveness of CI build predictions depends on the context, such as project characteristics and the individual experience levels of developers. This observation underscores the potential benefits of tailoring the level of automation and the nature of interactions between the build prediction technique and its users based on these contexts. It suggests that a one-size-fits-all approach may not be optimal and that customization could enhance the build prediction’s effectiveness and user satisfaction. It is worth conducting further research involving in-depth case studies or comparative analyses across diverse contexts to understand the contextual factors affecting the utilization of CI builds prediction in practice.
Enhancement of Build Prediction Explainability: Our observation from RQ4 pinpoints the importance of explainability in CI build prediction. We found that developers are more likely to trust and effectively utilize these techniques when they understand the rationale behind predictions. Although we give a glance into the explanation of the CI build prediction, there’s substantial room for advancement. Future research can focus on creating intuitive methods to convey complex model decisions such as interactive visualizations or clear narrative explanations. Studies could also explore the impact of different explanation types on developer trust and model utilization, aiming to tailor explainability features to developers’ specific needs and preferences. Enhancing explainability plays an important role in fostering relationships between developers and CI build prediction.
5.2. Implications for Practitioners